Mr. James gives a map to a first student studying Fisheries and Aquaculture so that she gets to the new fish ponds. The map says, ‘start from the central administration car park, go 125m straight north, then 200m at 45o north of east, then 150m straight east, and finally, 30m at 30o east of south’. How far from the central administration car park and in what direction are the fish ponds?


1
Expert's answer
2021-03-08T07:33:00-0500
"s_x=200\cos{45}+150+30\sin{30}=306.42\ m\\
s_y=125+200\sin{45}-30\cos{30}=240.44\ m"

"s=\sqrt{306.42^2+240.44^2}=389.5\ m\\
\theta=\arctan{\frac{240.44}{306.42}}=38.1\degree \text{ N of E}"